OpenStudy (anonymous):

Verify the identity. (cot x)/(1+cos x) = (csc x - 1)/(cot x)

OpenStudy (dan815):

show that cot^2x/(1+cosx)=cscx-1 csc=1/sin cot=cos/sin

OpenStudy (anonymous):

Can someone break this down further? I don't understand.